OpenStudy (anonymous):

When the integral from -2 to 5 has h(x)dx=-11. How do you solve the integral from 0 to 7 has h(x-2)dx=?

OpenStudy (anonymous):

You need to do a u substitution. you have this right now:\[\int\limits_{-2}^{5}h(x)dx = -11\]and the question is:\[\int\limits_{0}^{7}h(x-2)dx=?\]So start with:\[\int\limits_{0}^{7}h(x-2)dx\]Let u = x-2, then du=dx, and your lower limit changes to:\[0-2=-2\]while your upper limit becomes:\[7-2=5\]So with your substitution in place, you see that:\[\int\limits_{0}^{7}h(x-2)dx=\int\limits_{-2}^{5}h(u)du\]

OpenStudy (anonymous):

Thank you!!

OpenStudy (anonymous):

Wow, ok so that means the answer IS -11.

OpenStudy (anonymous):

yes, thats correct :)

OpenStudy (anonymous):

I'm in it for the "how to." This tells me how to solve the last fourth of my homework.